Charley pays $4.00 to play a game in which he flips two coins. If both coins land on heads, he wins $10.00. If the coins land any other way, he will loose the $4.00 he paid to play. What is the expected value to Charley for this game?
Answer by Boreal(15235)

E(X)=(1/4)*10-(3/4)*4=$2.50-$3.00=-$0.50.
BUT
he pays to play the game, so if he wins, he gets $10 minus $4 or $6.
E(X)=(1/4)*6-(3/4)*4=$1.50-$3.00=-$1.50
It needs to be clear whether he gets back the money he paid to play. I suspect he doesn't, so the expected value is -$1.50.